OpenStudy (mlgpro):

Which data set has the largest mode? 0, 0, 0, 1, 2, 2 0, 0, 1, 1, 1, 2 0, 1, 2, 3, 4, 5 0, 0, 1, 2, 3, 4

OpenStudy (anonymous):

its between the first and second choices

OpenStudy (mlgpro):

no those are the same ?

OpenStudy (anonymous):

by any chance would there be an option to choose both answers?

OpenStudy (mlgpro):

no those are the only choices

OpenStudy (mlgpro):

A=5/6 ? B=5/6 ? C=12/6 ? D=10/6 ?